Section 10

Partial Differential Equations

In case of several authors, Invited Speakers are marked with a ^{∗}.

Fabrice Bethuel: Vortices in Ginzburg-Landau Equations . . . . . . . . . III 11

Frédéric Hélein: Phenomena of Compensation and Estimates for Partial Differential Equations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 21

Robert R. Jensen: Viscosity Solutions of Elliptic Partial Differential Equations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 31

Hans Lindblad: Minimal Regularity Solutions of Nonlinear Wave Equations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 39

M. Machedon: Fourier Analysis of Null Forms and Non-linear Wave Equations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 49

Frank Merle: Blow-up Phenomena for Critical Nonlinear Schrödinger and Zakharov Equations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 57

Gustavo Ponce: On Nonlinear Dispersive Equations . . . . . . . . . . . . . . . III 67

Gunther Uhlmann: Inverse Boundary Value Problems for Partial Differential Equations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . III 77

D. Yafaev: Scattering Theory: Some Old and New Problems . . . . . . . . III 87

Documenta Mathematica · Extra Volume ICM 1998 · III ·

10

Documenta Mathematica · Extra Volume ICM 1998 · III ·